OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (2, -2) and (-1, 10) please show work!

OpenStudy (anonymous):

i don't understand what to do

OpenStudy (anonymous):

y=-4x+(10/4)

OpenStudy (anonymous):

how did you get y=-4

OpenStudy (anonymous):

to find m. which is slope of the form Y=MX+B \[m=\frac{y _{2}-y _{1}}{x _{2}-x _{1}}\]

OpenStudy (anonymous):

over three, down twelve same as over one down 4

OpenStudy (anonymous):

\[m=\frac{10-(-2)}{-1-2}\] \[m=\frac{12}{-3}\] \[m=-4\]

OpenStudy (anonymous):

now to find b. Just plug into the equation \[y=-4x+b\] the x and y and solve for b

OpenStudy (anonymous):

from -1 to 2 is three steps to the right on horizontally. from 10 to -2 is twelve steps down. over 3 down 12. same as over 1, down 4. that explains the -4 as the slope. (- because down)

OpenStudy (amistre64):

i tend to get my xs and ys misplaced so I perfer stacking and subtracting .... (2, -2) -(-1, 10) --------- 3,-12; slope = y/x = -12/3 = -4

OpenStudy (amistre64):

y = -4x -(-4.2) -2 y = -4x +8-2 y = -4x +6

OpenStudy (anonymous):

keebler you got this?

OpenStudy (anonymous):

yes